# Variance-reduction for Variational Inequality Problems with Bregman Distance Function

### v2

@@ 1 Introduction
One of the main structures commonly arise in a large-scale setting is when the operator $F$ consists of a large (finite) number of components, i.e., $F=\frac{1}{n}\sum_{i=1}^{n}F_{i}$. In this setting, when $n$ is large, evaluation of the operator $F$ (which may correspond to first-order information) can be prohibitively expensive. While stochastic methods have been extensively studied for a more general setting when the operator $F$ is represented by the expectation of a random variable, there is a significant gap in the convergence rate guarantee compared to deterministic methods. Consequently, variance reduction techniques have been developed to fill this gap to improve the complexity of deterministic methods. While variance reduction techniques have been successfully applied to minimization problems [4, 23], their naive extension to VI problems does not improve the complexity [3]. On the other hand, proximal-gradient-based methods with Bregman divergence have been shown to enhance algorithmic flexibility across various contexts, resulting in notable computational efficiencies in specific scenarios [46, 1]. The integration of Bregman distance with variance reduction for saddle point and VI problems poses an additional layer of complexity to algorithmic analysis. These challenges have been addressed by introducing retraction steps, a novel approach first implemented in the context of convex-concave saddle point problems [49] and later in [3] for monotone VIs. While prior methods rely on SVRG-based techniques, our approach introduces a novel variance-reduction algorithm inspired by SPIDER, offering a unified framework for solving both monotone and non-monotone VIs. This novel integration prompts the following research question:
@@ 1.1 Related work
In the past few decades, there have been various methods developed for solving VIs with convergence rate guarantee, such as Extragradient (EG)/Mirror-Prox (MP) [22, 29], forward-backward-forward (FBF) [43], dual extrapolation [31], or reflected gradient/forward-reflected-backward (FoRB) [27]. These methods typically demonstrate a complexity of $\mathcal{O}(1/\epsilon)$ when dealing with monotone VIs in deterministic settings. The significance of stochastic VIs has grown, particularly in scenarios involving uncertainty, where methods such as Sample Average Approximation (SAA) and Stochastic Approximation (SA) come into play. SAA involves approximating the expected value of stochastic mapping by averaging over a large number of samples, while SA employs a (mini-batch) sample at each iteration. Notably, the stochastic variants of these methods are characterized by a complexity of $\mathcal{O}(1/\epsilon^{2})$ when addressing monotone VIs [30, 22].
In optimization with finite-sum objectives, variance reduction techniques, pioneered by classical works such as [39, 21, 15], aim to enhance deterministic methods. They achieve this by offering unbiased estimators of gradients while minimizing the variance of the error in gradient estimation. Recently, there has been a growing interest in developing variance-reduced techniques for saddle point and VI problems which we discuss next.
@@
$\mathcal{O}(n+\tfrac{\sqrt{n}}{\epsilon})$
This paper $\mathcal{O}(n+\tfrac{\sqrt{n}}{\epsilon})$
@@
Finite-sum This Paper ✓(Lip.) $\mathcal{O}(n+\tfrac{1}{\epsilon^{2}})$
@@ 1.3 Application
Let $x=[u^{T}\lambda]^{T}$, $L_{i}(x,y)=ny_{i}\ell_{i}(x)-\lambda(\frac{1}{2}V_{i}(y_{i},\tfrac{1}{n}\mathbf{1}_{n})-\tfrac{\rho}{n})$, $h_{1}(x)=\mathbb{I}_{X\times\mathbb{R}_{+}}(u,\lambda)$ and $h_{2}(y)=\mathbb{I}_{\Delta_{n}}(y)$, then the previous equation will be a special case of SP problem which can be formulated as a VI by setting $F(z)=F(x,y)=\begin{bmatrix}\nabla_{x}L(x,y)\\ -\nabla_{y}L(x,y)\end{bmatrix}$ and $h(z)=h_{1}(x)+h_{2}(y)$.
Von Neumann’s Ratio Game: Consider a two-player game where the payoff function is of the form ${\langle x,Ry\rangle}/{\langle x,Sy\rangle}$, where $x,y$ are mixed-strategy vectors and $R\in\mathbb{R}^{n\times m}$, $S\in\mathbb{R}_{+}^{n\times n}$ are matrices. Such a payoff function arises in stochastic games, economic models of an expanding economy, and some nonzero-sum game formulations [32]. The goal is then to solve the following min-max problem
@@
where $X\subseteq\mathbb{R}^{n}$ is a convex and compact set, and we assume that $\langle x,Sy\rangle>0$, for all $(x,y)\in X\times\Delta_{m}$. Note that problem (4) can be formulated as (1) which has a solution [32]. Moreover, it has been shown that this problem satisfies the weak MVI condition [12].
@@ 2 Preliminaries
In this section, we provide some fundamental definitions and state the lemmas that we need for convergence analysis. Throughout the paper, $\|\cdot\|_{2}$ denotes the Euclidean norm and $\mathbb{I}_{X}(\cdot)$ denotes the indicator function of set $X$.
@@
Let $(\mathcal{U},\|\cdot\|_{\mathcal{U}})$ be a finite-dimensional normed vector space with the dual space $(\mathcal{U}^{*},\|\cdot\|_{\mathcal{U}^{*}})$, $f:\mathcal{U}\to\mathbb{R}\cup\{+\infty\}$ be a closed convex function, ${U}\subset\mathcal{U}$ is a closed convex set, and $\phi:\mathcal{U}\to\mathbb{R}$ be a distance-generating function which is continuously differentiable on an open set containing dom $\textbf{dom }f$ and is $l$-strongly convex with respect to .$\|.\|_{\mathcal{U}}$, and $\mathbf{D}_{U}:U\times(U\cap\textbf{dom }f)\to\mathbb{R}$ be a Bregman distance function associated with $\phi$. Then, the following result holds:
a) Given $\bar{x}\in U\cap\textbf{dom }f$, $s\in\mathcal{U}$, and $t>0$, let $x^{+}=\arg\min_{x\in U}f(x)-\langle s,x\rangle+t\mathbf{D}_{U}(x,\bar{x})$. Then, for all $x\in U$, the following inequality holds:
@@
b) For all $x\in U$ and $y,z\in U\cap\textbf{dom }f$, $\langle\nabla\phi(z)-\nabla\phi({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}y}),x-z\rangle=\mathbf{D}_{U}(x,y)-\mathbf{D}_{U}(x,z)-\mathbf{D}_{U}(z,y)$.
c) Given the update of $x^{+}$ in (a), for all $x\in U$ the following inequality holds:
@@
d) Assuming $\phi(.)$ is a closed function, then $\nabla\phi^{*}(\nabla\phi(x))=x$, for all $x\in\textbf{dom }\nabla\phi\subset\mathcal{U}$; and $\nabla\phi(\nabla\phi^{*}(y))=y$, for all $y\in\textbf{dom }\nabla\phi^{*}\subset\mathcal{U}^{*}$.
@@
For proofs see [44] and [5]. ∎
@@
$\tfrac{1}{2}\left\|x_{k+1}-\tilde{x}_{k}\right\|_{\mathcal{U}}^{2}\leq D^{k}_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}U}(x_{k+1}),$ (7a)
@@
$\left\|x_{k+1}-\tilde{x}_{k}\right\|_{\mathcal{U}}^{2}$ $=\left\|x_{k+1}-\frac{1}{q}\sum_{j=(i_{k}-1)q+1}^{i_{k}q}x_{j}\right\|_{\mathcal{U}}^{2}$
$=\left\|\frac{1}{q}\sum_{j=(i_{k}-1)q+1}^{i_{k}q}(x_{k+1}-x_{j})\right\|_{\mathcal{U}}^{2}$
@@ 3 Proposed method
Moreover, $h$ is possibly nonsmooth whose Bregman proximal operator, i.e., $\mathop{\rm argmin}_{x\in X}\break\{h(x)+\langle v,{x}\rangle+\tfrac{1}{\sigma}\mathbf{D}_{X}(x,\bar{x})\}$ for some $\bar{x},v\in\mathcal{X}$ and $\sigma>0$, can be computed efficiently.
@@
1:  Input: ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x}}_{0}\in{\mathcal{X}}$ , $S\in\{1,\ldots,n\}$, $q\in\{1,\ldots,n\}$, and $\beta\in[0,1]$
@@
For any $k\geq 0$, let $r_{k}\triangleq F_{S}(x_{k})-(1-\beta)F_{S}({x}_{k-1})-\beta F_{S}(\tilde{x}_{k-1})$, $\bar{r}_{k}\triangleq F(x_{k})-(1-\beta)F({x}_{k-1})-\beta F(\tilde{x}_{k-1})$, and $\delta_{k}\triangleq v_{k}-((1-\beta)F(x_{k})+\beta F(\tilde{x}_{k}))$.
@@
Let $\{x_{k}\}_{k\geq 0}$ be the sequence generated by Algorithm 1. We some define auxiliary sequence $\{w_{k}^{x},w_{k}^{r}\}\subset\mathcal{X}$, for $k\geq 0$ as follow:
@@
$\langle{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\delta_{k}},x-x_{k+1}\rangle$ $\leq\tfrac{1}{\eta_{k}^{1}}(\mathbf{D}_{X}(x,w_{k}^{x})-\mathbf{D}_{X}(x,w_{k+1}^{x}))+\left\langle\delta_{k},w_{k}^{x}-w_{k+1}^{x}+x_{k}-x_{k+1}\right\rangle$
$\quad-\tfrac{1}{\eta_{k}^{1}}\mathbf{D}_{X}(w_{k+1}^{x},w_{k}^{x})+\langle{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\delta_{k}},w_{k}^{x}-x_{k}\rangle,$ (10a)
@@
Next, recall that ${\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}r}_{k}=F_{S}(x_{k})-(1-\beta)F_{S}({x}_{k-1})-\beta F_{S}(\tilde{x}_{k-1})$ and $\bar{r}_{k}=F(x_{k})-(1-\beta)F({x}_{k-1})-\beta F(\tilde{x}_{k-1})$. It is easy to verify that $\mathbb{E}[r_{k}-\bar{r}_{k}]=0$. Define $\tilde{e}_{k}\triangleq F_{S}(x_{k})-F_{S}(\tilde{x}_{k-1})+F(\tilde{x}_{k-1})-F(x_{k})$, then using the triangle inequality and Assumption 1, one can easily demonstrate that
@@
$\quad+\tfrac{\gamma^{y}_{k}}{\sigma_{k}}\langle\nabla\psi_{\mathcal{X}}(x_{k+1})-s_{k},x-x_{k+1}\rangle$
$=\langle v_{k}+\theta_{k}r_{k},x-x_{k+1}\rangle+\tfrac{1-\gamma_{k}}{\sigma_{k}}\langle\nabla\psi_{\mathcal{X}}(x_{k+1})-\nabla\psi_{\mathcal{Y}}(x_{k}),x-x_{k+1}\rangle$
@@
Where in the first equality we used Lemma 1 part(d) and updated the rule of $\hat{x}_{k}$ in Algorithm 1. Then by using the generalized three-point property of Bregman distance in Lemma 1 part (b) twice; we will have:

### v3

@@ 1 Introduction
One of the main structures commonly arise in a large-scale setting is when the operator $F$ consists of a large (finite) number of components, i.e., $F=\frac{1}{n}\sum_{i=1}^{n}F_{i}$. In this setting, when $n$ is large, evaluation of the operator $F$ (which may correspond to first-order information) can be prohibitively expensive. While stochastic methods have been extensively studied for a more general setting when the operator $F$ is represented by the expectation of a random variable, there is a significant gap in the convergence rate guarantee compared to deterministic methods. Consequently, variance reduction techniques have been developed to fill this gap to improve the complexity of deterministic methods. While variance reduction techniques have been successfully applied to minimization problems [4, 25], their naive extension to VI problems does not improve the complexity [3]. On the other hand, proximal-gradient-based methods with Bregman divergence have been shown to enhance algorithmic flexibility across various contexts, resulting in notable computational efficiencies in specific scenarios [53, 1]. The integration of Bregman distance with variance reduction for saddle point and VI problems poses an additional layer of complexity to algorithmic analysis. These challenges have been addressed by introducing retraction steps, a novel approach first implemented in the context of convex-concave saddle point problems [56] and later in [3] for monotone VIs. While prior methods rely on SVRG-based techniques, our approach introduces a novel variance-reduction algorithm inspired by SPIDER (see ; [16]), offering a unified framework for solving both monotone and non-monotone VIs. This novel integration prompts the following research question:
@@ 1.1 Related work
In the past few decades, there have been various methods developed for solving VIs with convergence rate guarantee, such as Extragradient (EG)/Mirror-Prox (MP) [24, 33], forward-backward-forward (FBF) [50], dual extrapolation [35], or reflected gradient/forward-reflected-backward (FoRB) [30]. More specifically, FBF improves the classical forward-backward scheme by incorporating an extra forward step, and relaxing the need for cocoercivity assumption. FoRB, on the other hand, simplifies the update rule by using only one forward and one backward step while still maintaining convergence under monotonicity, making it suitable for problems where full extragradient steps are computationally expensive.While these methods differ in their use of operator calls and projections, they typically demonstrate a complexity of $\mathcal{O}(1/\epsilon)$ when dealing with monotone VIs in deterministic settings. The significance of stochastic VIs has grown, particularly in scenarios involving uncertainty, where methods such as Sample Average Approximation (SAA) and Stochastic Approximation (SA) come into play. SAA involves approximating the expected value of stochastic mapping by averaging over a large number of samples, while SA employs a (mini-batch) sample at each iteration. Notably, the stochastic variants of these methods are characterized by a complexity of $\mathcal{O}(1/\epsilon^{2})$ when addressing monotone VIs [34, 24].
In the optimization literature with a finite-sum objective function, variance reduction techniques, pioneered by classical works such as SVRG [23], SAGA [13], SARAH [37], and SPIDER [16], aim to improve deterministic methods. They achieve this by offering unbiased estimators of gradients while minimizing the variance of the error in gradient estimation. These algorithms obtain faster convergence rates–often linear under strong convexity–by using recursive update rules or control variates to refine noisy gradient estimates based on previously computed information, while preserving the low per-iteration computational cost characteristic of stochastic gradient methods. Recently, there has been a growing interest in developing variance-reduced techniques for saddle point and VI problems, which we discuss next.
@@
$\mathcal{O}(n+\frac{\sqrt{n}}{{\epsilon}})$
This paper $\mathcal{O}(n+\frac{\sqrt{n}}{{\epsilon}})$
@@
Finite-sum This Paper ✓(Lip.) $\mathcal{O}(n+\frac{1}{\epsilon^{2}})$
@@ 1.3 Application
Let $x=[u^{T}\lambda]^{T}$, $L_{i}(x,y)=ny_{i}\ell_{i}(x)-\lambda(\frac{1}{2}V_{i}(y_{i},\tfrac{1}{n}\mathbf{1}_{n})-\tfrac{\rho}{n})$, $h_{1}(x)=\mathbb{I}_{X\times\mathbb{R}_{+}}(u,\lambda)$ and $h_{2}(y)=\mathbb{I}_{\Delta_{n}}(y)$, where $\mathbb{I}_{X}(\cdot)$ denotes the indicator function of set $X$, then the previous equation will be a special case of SP problem which can be formulated as a VI by setting $F(z)=F(x,y)=\begin{bmatrix}\nabla_{x}L(x,y)\\ -\nabla_{y}L(x,y)\end{bmatrix}$ and $h(z)=h_{1}(x)+h_{2}(y)$.
Von Neumann’s Ratio Game: Consider a two-player game where the payoff function is of the form ${\langle x,Ry\rangle}/{\langle x,Sy\rangle}$, where $x,y$ are mixed-strategy vectors and $R\in\mathbb{R}^{n\times m}$, $S\in\mathbb{R}_{+}^{n\times n}$ are matrices. Such a payoff function arises in stochastic games, economic models of an expanding economy, and some nonzero-sum game formulations [36]. The goal is then to solve the following minimax problem
@@
where $X\subseteq\mathbb{R}^{n}$ is a convex and compact set, and we assume that $\langle x,Sy\rangle>0$, for all $(x,y)\in X\times\Delta_{m}$. Note that problem (4) can be formulated as (1) which has a solution [36]. Moreover, this problem has been shown to satisfy the weak MVI condition [12].
@@ 2 Preliminaries
In this section, we provide some fundamental definitions and state the lemmas that we need for the convergence analysis. Throughout the paper, $\|\cdot\|_{2}$ denotes the Euclidean norm.
@@
Let $(\mathcal{U},\|\cdot\|_{\mathcal{U}})$ be a finite-dimensional normed vector space with the dual space $(\mathcal{U}^{*},\|\cdot\|_{\mathcal{U}^{*}})$, $f:\mathcal{U}\to\mathbb{R}\cup\{+\infty\}$ be a closed convex function, ${U}\subset\mathcal{U}$ is a closed convex set, and $\phi:\mathcal{U}\to\mathbb{R}$ be a distance-generating function which is continuously differentiable on an open set containing dom $\textbf{dom }f$ and is $1$-strongly convex with respect to $\|\cdot\|_{\mathcal{U}}$, and $\mathbf{D}_{U}:U\times(U\cap\textbf{dom }f)\to\mathbb{R}$ be a Bregman distance function associated with $\phi$. Then, the following result holds:
(a) For all $x\in U$ and $y,z\in U\cap\textbf{dom }f$, $\langle\nabla\phi(z)-\nabla\phi({\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}y}),x-z\rangle=\mathbf{D}_{U}(x,y)-\mathbf{D}_{U}(x,z)-\mathbf{D}_{U}(z,y)$.
(b) Given $\bar{x}\in U\cap\textbf{dom }f$, $s\in\mathcal{U}$, and $t>0$, let $x^{+}=\arg\min_{x\in U}f(x)-\langle s,x\rangle+t\mathbf{D}_{U}(x,\bar{x})$. Then, for all $x\in U$, the following inequality holds:
@@
(c) Given the update of $x^{+}$ in (a), for all $x\in U$ the following inequality holds:
@@
(d) Assuming $\phi(\cdot)$ is a closed function, then $\nabla\phi^{*}(\nabla\phi(x))=x$, for all $x\in\textbf{dom }\nabla\phi\subset\mathcal{U}$; and $\nabla\phi(\nabla\phi^{*}(y))=y$, for all $y\in\textbf{dom }\nabla\phi^{*}\subset\mathcal{U}^{*}$.
@@
Part $(a)$ follows from the definition of the Bregman distance function. To prove part $(b)$, one can apply [51, Property 1] combined with part $(a)$. Part $(c)$ follows by adding $\left\langle s,\bar{x}-x\right\rangle$ to both sides of (5) and using Young’s inequality as follows:
$f(x)+\left\langle s,\bar{x}-x\right\rangle$ $\geq f(x^{+})+\langle s,\bar{x}-x^{+}\rangle+t\langle\nabla\phi(\bar{x})-\nabla\phi(x^{+}),x-x^{+}\rangle$
$\geq f(x^{+})-\frac{t}{2}\|\bar{x}-x^{+}\|^{2}_{\mathcal{U}}-\frac{1}{2t}\|s\|^{2}_{\mathcal{U}^{*}}+t\langle\nabla\phi(\bar{x})-\nabla\phi(x^{+}),x-x^{+}\rangle$
$\geq f(x^{+})-\frac{1}{2t}\|s\|^{2}_{\mathcal{U}^{*}}-t\left(\mathbf{D}_{U}(x,\bar{x})-\mathbf{D}_{U}(x,x^{+})\right),$
where in the last inequality, we used the 1-strong convexity of the Bregman distance and part $(a)$. Finally, part $(d)$ follows from [44, Theorem 26.5]. ∎
@@
${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\tfrac{1}{2}}}\left\|x_{k+1}-\tilde{x}_{k}\right\|_{\mathcal{U}}^{2}\leq D^{k}_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}U}(x_{k+1}),$ (7a)
@@
$\left\|x_{k+1}-\tilde{x}_{k}\right\|_{\mathcal{U}}^{2}$ $=\Big{\|}{x_{k+1}-\frac{1}{q}\sum_{j=(i_{k}-1)q+1}^{i_{k}q}x_{j}}\Big{\|}_{\mathcal{U}}^{2}$
$=\Big{\|}{\frac{1}{q}\sum_{j=(i_{k}-1)q+1}^{i_{k}q}(x_{k+1}-x_{j})}\Big{\|}_{\mathcal{U}}^{2}$
@@ 3 Algorithmic Framework
Moreover, $h$ is possibly nonsmooth whose Bregman proximal operator, i.e., $\mathop{\rm argmin}_{x\in X}\break\{h(x)+\langle v,{x}\rangle+\tfrac{1}{\sigma}\mathbf{D}_{X}(x,\bar{x})\}$ for some $\bar{x},v\in\mathcal{X}$ and $\sigma>0$, can be computed efficiently. The primary motivation for introducing Bregman distances in our algorithm design is the ability to better capture the geometry of the feasible set, especially in structured, non-Euclidean settings. For example, the effectiveness of using the negative entropy Bregman distance for projection onto the simplex set has been studied in [26, 31, 46].
@@
1:  Input: ${{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}x}}_{0}\in{\mathcal{X}}$ , $S\in\{1,\ldots,n\}$, $q\in\{1,\ldots,n\}$, $\beta\in[0,1]$, $\theta_{k}=1,\gamma_{k}\in(0,1)$ and $\sigma_{k}>0$
@@
For any $k\geq 0$, let $\delta_{k}\triangleq v_{k}-((1-\beta)F(x_{k})+\beta F(\tilde{x}_{k}))$, $\bar{r}_{k}\triangleq F(x_{k})-(1-\beta)F({x}_{k-1})-\beta F(\tilde{x}_{k-1})$, and
$r_{k}\triangleq\begin{cases}F_{S}(x_{k})-(1-\beta)F_{S}({x}_{k-1})-\beta F_{S}(\tilde{x}_{k-1})&if\quad\text{mod}(k,q)\neq 0\\ \bar{r}_{k}&o.w.\end{cases}$
@@
Let $\{x_{k}\}_{k\geq 0}$ be the sequence generated by Algorithm 1. We define some auxiliary sequence $\{w_{k}^{x},w_{k}^{r}\}\subset\mathcal{X}$, for $k\geq 0$ as follow:
@@
$\langle{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\delta_{k}},x-x_{k+1}\rangle$ $\leq\tfrac{1}{\eta_{k}^{1}}(\mathbf{D}_{X}(x,w_{k}^{x})-\mathbf{D}_{X}(x,w_{k+1}^{x}))+{{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\left\langle\delta_{k},w_{k}^{x}-w_{k+1}^{x}+x_{k}-x_{k+1}\right\rangle}}$
$\quad-\tfrac{1}{\eta_{k}^{1}}\mathbf{D}_{X}(w_{k+1}^{x},{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}w}_{k}^{x})+\langle{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\delta_{k}},w_{k}^{x}-x_{k}\rangle,$ (10a)
@@
Next, recall that for any $k\geq 0$, $\bar{r}_{k}=F(x_{k})-(1-\beta)F({x}_{k-1})-\beta F(\tilde{x}_{k-1})$, moreover, $r_{k}=F_{S}(x_{k})-(1-\beta)F_{S}({x}_{k-1})-\beta F_{S}(\tilde{x}_{k-1})$ if $\text{mod}(k,q)\neq 0$ and $\bar{r}_{k}$ otherwise. It is easy to verify that $\mathbb{E}[r_{k}-\bar{r}_{k}]=0$ for $k\geq 0$. Define $\tilde{e}_{k}\triangleq F_{S}(x_{k})-F_{S}(\tilde{x}_{k-1})+F(\tilde{x}_{k-1})-F(x_{k})$, then using the triangle inequality and Assumption 1, one can easily demonstrate that
@@
$\quad+\tfrac{\gamma_{k}}{\sigma_{k}}\langle\nabla\psi_{\mathcal{X}}(x_{k+1})-s_{k},x-x_{k+1}\rangle$
@@
Where in the first equality we used Lemma 1 part(d) and updated the rule of $\hat{x}_{k}$ in Algorithm 1. Then by using the generalized three-point property of Bregman distance in Lemma 1 part (a) twice; we will have: